# Perverse filtrations and Fourier transforms

### v2

@@ 0.2. Dualizable abelian fibrations
Let $C\to B$ be a flat family of integral projective curves which admits a section. If the associated compactified Jacobian fibration $\pi:\overline{J}_{C}\to B$ satisfies that
@@
    every curve in the family has at worst planar singularities, and
@@ 0.3.2. The $P=W$ conjecture
[…]
@@ 0.3.3. Enumerative geometry of local ${\mathbb{P}}^{2}$
Fix coprime integers $r,\chi$ with $r\geq 3$. We consider the moduli $M_{r,\chi}$ of $1$-dimensional stable sheaves $F$ on ${\mathbb{P}}^{2}$ with $[\mathrm{supp}(F)]=rH,\,\chi(F)=\chi$. Here $H$ is the hyperplane class. The moduli space $M_{r,\chi}$ admits a proper Hilbert–Chow morphism
@@
given by integrating the normalized $\mathrm{ch}_{k+1}$ of a universal family over $H^{j}\in H^{2j}({\mathbb{P}}^{2},{\mathbb{Q}})$; see Section 5.3.2. Moreover, it was shown in [49] that the first $73r-$ 7 tautological classes
@@ 1.2. The Beauville decomposition
Let $\pi:A\to B$ be an abelian scheme of relative dimension $g$ over a nonsingular base variety $B$. The “multiplication by $N$” map $[N]:A\to A$ induces a canonical decomposition of the cohomology
@@
is an eigenspace of the pullback map $[N]^{*}:H^{*}(A,{\mathbb{Q}})\to H^{*}(A,{\mathbb{Q}})$. The decomposition (7) is motivic (i.e. induced by projectors) and provides a splitting of the Leray filtration [^2] Here $L_{k}H^{d}(A,{\mathbb{Q}})=L^{d-k}H^{d}(A,{\mathbb{Q}})$ with respect to the usual convention for the Leray filtration. $L_{\bullet}H^{*}(A,{\mathbb{Q}})$ associated with the fibration $\pi:A\to B$, where
@@
In fact, since for $\alpha,\beta\in H^{*}(A,{\mathbb{Q}})$ we have $[N]^{*}(\alpha\cup\beta)=[N]^{*}\alpha\cup[N]^{*}\beta$, the multiplicativity follows simply by comparing the eigenvalues. For our purpose, we would like to extend (7) to abelian fibrations with singular fibers. We give another description of (7) using duality and Fourier transforms.
@@ 2.2.3. Multiplicative structure
[…]
@@
corresponding to the small relative diagonal $[\Delta^{\mathrm{sm}}_{X/B}]\in\mathrm{Corr}^{0}_{B}(X,X;X)$. It specializes to the sheaf-theoretic cup-product in $D^{b}_{c}(B)$,
@@
by restricting to the closed point $b\in B$ [^7] In the Chow realization, the cup-product specializes to the intersection product on $\mathrm{CH}^{*}(X)$.
@@ 2.4. Statement of the main theorem
(26) $\mathfrak{F}^{-1}$ $:={\mathrm{td}}(-(\pi\times_{B}\pi^{\vee})^{*}T_{B})\cap\tau({\mathcal{P}}^{-1})$
@@ 2.5.1. Decomposition
$[\Delta_{M/B}]={\mathfrak{F}}\circ{\mathfrak{F}}^{-1}=\sum_{i=0}^{2g}{\mathfrak{F}}_{i}\circ{\mathfrak{F}}^{-1}_{2g-i}={\mathfrak{p}}_{k}+{\mathfrak{q}}_{k+1}\in\mathrm{Corr}_{B}^{0}(M,M),$
@@
This shows that ${\mathfrak{p}}_{k}$ and ${\mathfrak{q}}_{k+1}$ are orthogonal projectors adding up to $[\Delta_{M/B}]$, which proves the motivic decomposition (29).
@@ 2.5.3. Multiplicativity
[…]
@@
[…]
@@
[…]
@@ 3.1. Overview
In this section, we treat the geometry of (degree 00) compactified Jacobian fibrations; our main purpose is to prove Theorem 0.2.
Throughout this section, we let $C\to B$ be a flat family of integral projective curves of arithmetic genus $g$ with planar singularities over an irreducible base $B$. We further assume that $C\to B$ admits a section whose image is contained in the smooth locus of the fibers. We denote by $\pi:\overline{J}_{C}\to B$ the compactified Jacobian fibration, and assume that the total space $\overline{J}_{C}$ is nonsingular. In particular, $B$ is nonsingular and $\pi:\overline{J}_{C}\to B$ is flat with integral fibers of dimension $g$.
@@ 3.5.3. Proof of (FV) for $\pi:\overline{J}_{C}\to B$
[…]
The dimension bound of Corollary 3.5 now implies that both ends of (61) vanish in codimensions $<g$. In other words, we have for $k<2g$,
@@ 3.6. Variants
Recall from [5, 6] that Arinkin’s normalized Poincaré sheaf ${\mathcal{P}}$ on $\overline{J}_{C}\times_{B}\overline{J}_{C}$ is constructed from the normalized universal family ${\mathcal{F}}$ on $C\times_{B}\overline{J}_{C}$ of degree 00 rank $1$ torsion-free sheaves. Here the normalization of ${\mathcal{F}}$ is characterized by the trivializations along the sections
@@
Using the section of $C\to B$, we can identify the degree $d$ compactified Jacobian $\overline{J}^{d}_{C}$ with the (degree 00) compactified Jacobian $\overline{J}_{C}$. A universal family of rank $1$ degree $d$ torsion-free sheaves on $C\times_{B}\overline{J}^{d}_{C}=C\times_{B}\overline{J}_{C}$ is given by
@@ 4.2. Universal families and gerbes
This is representable by a Deligne–Mumford stack. Finally it is a $\mu_{r}$-gerbe over $\overline{J}^{d}_{C}$ since the perfect complex ${\mathcal{R}}^{d}_{\overline{\mathfrak{J}}^{d}_{C}}$ has rank $r$ (which can be seen directly from nonsingular fibers). ∎
@@
It suffices to show that their cones in the derived categories are 00; this is a property which can be checked étale locally over the base $B$. We postpone the étale local calculations to Corollary 4.5 below. ∎
@@
They are independent and do not have any nontrivial relation. Due to the existence of the section, the relative compactified Jacobian $\overline{J}^{d}_{C}$ are identified for any choices of $d$. As before, we denote by $\overline{J}_{C}$ the degree 00 compactified Jacobian, and we want to compare ${\mathcal{F}}^{d}$ on $C\times_{U}\overline{{\mathcal{J}}}^{d}_{C}$ (which is trivialized along the multisection $D$) with the normalized universal sheaf ${\mathcal{F}}$ on $C\times_{U}\overline{J}_{C}$ (which is trivialized along the section $U$).
@@
Note that $\overline{J}_{C}$ is a fine moduli space. Therefore, describing a morphism from $\overline{{\mathcal{J}}}^{d}_{C}$ is equivalent to constructing a family of rank $1$ torsion-free sheaves of degree 00 trivialized along the section $U\subset C$, for which we may use ${\mathcal{F}}^{d}$. More precisely, we define
@@
Here the second factor adjusts the degree from $d$ to 00, and the third factor is included to trivialize the universal sheaf along the section. This yields a morphism $\sigma_{d}:\overline{{\mathcal{J}}}^{d}_{C}\to\overline{J}_{C}$ satisfying
@@
The normalized Poincare sheaf ${\mathcal{P}}$ and its inverse ${\mathcal{P}}^{-1}$ are canonically defined on the target. The following is an immediate consequence of Proposition 4.4 as in Section 3.6 (using the argument of [1, Proposition 3.1(i)]). It provides an étale local description of the Poincaré sheaf ${\mathcal{P}}_{e,d}$ and its inverse ${\mathcal{P}}^{-1}_{e,d}$.
@@ 5.3.1. The locus of integral curves
$[\mathrm{supp}(F)]=rH,\quad\chi(F)=\chi.$
@@
$\mathrm{codim}_{{\mathbb{P}}H^{0}({\mathbb{P}}^{2},{\mathcal{O}}_{{\mathbb{P}}^{2}}(r))}({\mathbb{P}}H^{0}({\mathbb{P}}^{2},{\mathcal{O}}_{{\mathbb{P}}^{2}}(r))\backslash W)=r-1.$
@@
This is well-known: the component given by the closure of reducible curves with two components of bidegrees $(1,r-1)$ has the largest dimension among all the boundary components. ∎
@@
$i_{*}i^{!}{\mathbb{Q}}_{Z}\to{\mathbb{Q}}_{M_{r,\chi}}\to j_{*}j^{*}{\mathbb{Q}}_{M_{W}}\to i_{*}i^{!}{\mathbb{Q}}_{Z}[1]$
@@
(71) $h_{*}i_{*}i^{!}{\mathbb{Q}}_{Z}\to h_{*}{\mathbb{Q}}_{M_{r,\chi}}\xrightarrow{~{}~{}(*)~{}~{}}j_{*}h_{W*}{\mathbb{Q}}_{M_{W}}\to h_{*}i_{*}i^{!}{\mathbb{Q}}_{Z}[1].$
The filtered morphism $\mathrm{res}_{W}$ is induced by taking the global cohomology of ()(*)( ∗ ). By Lemma 5.2, we have
$i_{*}i^{!}{\mathbb{Q}}_{Z}\in D_{c}^{\geq 2(r-1)}(M_{r,\chi}),\quad i_{*}i^{!}{\mathbb{Q}}_{Z}\in{{}^{\mathfrak{p}}}D_{c}^{\geq 2(r-1)}(M_{r,\chi}).$
In particular, both the first and the last terms of (71) are concentrated in degrees $>2(r-2)$, and
${}^{\mathfrak{p}}{\mathcal{H}}^{i}(*):{{}^{\mathfrak{p}}}{\mathcal{H}}^{i}(h_{*}{\mathbb{Q}}_{M_{r,\chi}})\rightarrow{~{}~{}}^{\mathfrak{p}}{\mathcal{H}}^{i}(j_{*}h_{W*}{\mathbb{Q}}_{M_{W}})$
is an isomorphism for any $i\leq 2(r-2)$. The corollary follows. ∎

### v3

@@ 0.2. Dualizable abelian fibrations
Let $C\to B$ be a flat family of integral projective curves which admits a section through its smooth locus. Let $\pi:\overline{J}_{C}\to B$ be the associated compactified Jacobian fibration. If
@@
    every curve in the family $C\to B$ has at worst planar singularities, and
@@ 0.3.2. The $P=W$ conjecture
(2) $\text{``}P=W\text{''},\quad P_{k}H^{*}(M_{\mathrm{Dol}},{\mathbb{Q}})=W_{2k}H^{*}(M_{\mathrm{B}},{\mathbb{Q}}).$
@@ 0.3.3. Enumerative geometry of local ${\mathbb{P}}^{2}$
Fix coprime integers $r,\chi$ with $r\geq 3$. We consider the moduli $M_{r,\chi}$ of $1$-dimensional stable sheaves $F$ on ${\mathbb{P}}^{2}$ with $\mathrm{supp}(F)]=rH,\,\chi(F)=\chi$. Here $H$ is the hyperplane class. The moduli space $M_{r,\chi}$ admits a proper Hilbert–Chow morphism
@@
given by integrating the normalized $\mathrm{ch}_{k+1}$ of a universal family over $H^{j}\in H^{2j}({\mathbb{P}}^{2},{\mathbb{Q}})$; see Section 5.3.2. Moreover, it was shown in [49] that the first 3r73r-7 tautological classes
@@ 1.2. The Beauville decomposition
Let $\pi:A\to B$ be an abelian scheme of relative dimension $g$ over a nonsingular base variety $B$. The “multiplication by NN” map $N]:A\to A$ induces a canonical decomposition of the cohomology
@@
is an eigenspace of the pullback map $N]^{*}:H^{*}(A,{\mathbb{Q}})\to H^{*}(A,{\mathbb{Q}})$. The decomposition (7) is motivic (i.e. induced by projectors) and provides a splitting of the Leray filtration [^2] Here $L_{k}H^{d}(A,{\mathbb{Q}})=L^{d-k}H^{d}(A,{\mathbb{Q}})$ with respect to the usual convention for the Leray filtration. $L_{\bullet}H^{*}(A,{\mathbb{Q}})$ associated with the fibration $\pi:A\to B$, where
@@
In fact, since for $\alpha,\beta\in H^{*}(A,{\mathbb{Q}})$ we have $N]^{*}(\alpha\cup\beta)=[N]^{*}\alpha\cup[N]^{*}\beta$, the multiplicativity follows simply by comparing the eigenvalues. For our purpose, we would like to extend (7) to abelian fibrations with singular fibers. We give another description of (7) using duality and Fourier transforms.
@@ 2.2.3. Multiplicative structure
$\simeq{}$ $R\mathcal{H}\mathrm{om}\left(f_{1}^{*}f_{2*}{\mathbb{Q}}_{X_{2}}[j],f_{1}^{!}f_{3*}{\mathbb{Q}}_{X_{3}}[k]\right)$
$\textrm{($f_{2}$ is proper)}\simeq{}$ $R\mathcal{H}\mathrm{om}\left(p^{12}_{1*}{\mathbb{Q}}_{X_{1}\times_{B}X_{2}}[j],p^{13}_{1*}p^{13!}_{3}{\mathbb{Q}}_{X_{3}}[k]\right)$
$\textrm{($X_{3}$ is nonsingular)}\simeq{}$ $R\mathcal{H}\mathrm{om}\left(p^{12}_{1*}{\mathbb{Q}}_{X_{1}\times_{B}X_{2}}[j],p^{13}_{1*}\omega_{X_{1}\times_{B}X_{3}}[k-2\dim X_{3}]\right)$
$\textrm{($X_{1}$ is nonsingular)}\simeq{}$ $R\mathcal{H}\mathrm{om}\left(p^{12}_{1*}{\mathbb{Q}}_{X_{1}\times_{B}X_{2}}[j],p^{13}_{1*}p^{13!}_{1}{\mathbb{Q}}_{X_{1}}[k+2\dim X_{1}-2\dim X_{3}]\right)$
$\textrm{($p^{12}_{1}$ is proper)}\simeq{}$ $p^{12}_{1*}R\mathcal{H}\mathrm{om}\left({\mathbb{Q}}_{X_{1}\times_{B}X_{2}}[j],p^{12!}_{1}p^{13}_{1*}p^{13!}_{1}{\mathbb{Q}}_{X_{1}}[k+2\dim X_{1}-2\dim X_{3}]\right)$
$\simeq{}$ $p^{12}_{1*}\left(p^{12!}_{1}p^{13}_{1*}\right)p^{13!}_{1}{\mathbb{Q}}_{X_{1}}[k-j+2\dim X_{1}-2\dim X_{3}]$
$\simeq{}$ $p^{12}_{1*}\left(p^{123}_{12*}p^{123!}_{13}\right)p^{13!}_{1}{\mathbb{Q}}_{X_{1}}[k-j+2\dim X_{1}-2\dim X_{3}]$
$\simeq{}$ $p^{123}_{1*}p^{123!}_{1}{\mathbb{Q}}_{X_{1}}[k-j+2\dim X_{1}-2\dim X_{3}],$
@@
corresponding to the small relative diagonal $\Delta^{\mathrm{sm}}_{X/B}]\in\mathrm{Corr}^{0}_{B}(X,X;X)$. It specializes to the sheaf-theoretic cup-product in $D^{b}_{c}(B)$,
@@
by restricting to the closed point $b\in B$. [^7] In the Chow realization, the cup-product specializes to the intersection product on $\mathrm{CH}^{*}(X)$.
@@ 2.4. Statement of the main theorem
(26) $\mathfrak{F}^{-1}$ $={\mathrm{td}}(-(\pi\times_{B}\pi^{\vee})^{*}T_{B})\cap\tau({\mathcal{P}}^{-1})$
@@ 2.5.1. Decomposition
$\Delta_{M/B}]={\mathfrak{F}}\circ{\mathfrak{F}}^{-1}=\sum_{i=0}^{2g}{\mathfrak{F}}_{i}\circ{\mathfrak{F}}^{-1}_{2g-i}={\mathfrak{p}}_{k}+{\mathfrak{q}}_{k+1}\in\mathrm{Corr}_{B}^{0}(M,M),$
@@
This shows that ${\mathfrak{p}}_{k}$ and ${\mathfrak{q}}_{k+1}$ are orthogonal projectors adding up to $\Delta_{M/B}]$, which proves the motivic decomposition (29).
@@ 2.5.3. Multiplicativity
${\mathfrak{F}}\circ{\mathfrak{C}}\circ({\mathfrak{F}}^{-1}\times{\mathfrak{F}}^{-1})=[\Delta^{\mathrm{sm}}_{M/B}]\in\mathrm{Corr}_{B}^{0}(M,M;M).$
@@
${}={}$ $\left(\sum_{i_{3}\geq k+l+1}{\mathfrak{F}}_{i_{3}}\circ{\mathfrak{F}}^{-1}_{2g-i_{3}}\right)\circ\left(\sum_{j_{3}\geq k+l+1}{\mathfrak{F}}_{j_{3}}\right)\circ{\mathfrak{C}}$
@@
${}={}$ ${\mathfrak{q}}_{k+l+1}\circ\left(\sum_{j_{3}\geq k+l+1}{\mathfrak{F}}_{j_{3}}\right)\circ{\mathfrak{C}}\circ\left(\left(\left(\sum_{j_{1}\geq 2g-k}{\mathfrak{F}}^{-1}_{j_{1}}\right)\circ{\mathfrak{p}}_{k}\right)\times\left(\left(\sum_{j_{2}\geq 2g-l}{\mathfrak{F}}^{-1}_{j_{2}}\right)\circ{\mathfrak{p}}_{l}\right)\right).$
@@ 3.1. Overview
In this section, we treat the geometry of (degree 0) compactified Jacobian fibrations; our main purpose is to prove Theorem 0.2.
Throughout this section, we let $C\to B$ be a flat family of integral projective curves of arithmetic genus $g$ with planar singularities over an irreducible base $B$. We further assume that $C\to B$ admits a section with image contained in its smooth locus. We denote by $\pi:\overline{J}_{C}\to B$ the compactified Jacobian fibration, and assume that the total space $\overline{J}_{C}$ is nonsingular. In particular, $B$ is nonsingular and $\pi:\overline{J}_{C}\to B$ is flat with integral fibers of dimension $g$.
@@ 3.5.3. Proof of (FV) for $\pi:\overline{J}_{C}\to B$
$={}$ ${\mathrm{td}}(-q_{1}^{*}T_{\overline{J}_{C}}-\pi_{2}^{*}T_{B})\cap p_{13*}\left({\mathrm{td}}(-p_{2}^{*}T_{\overline{J}_{C}})\cap\delta^{!}(\tau(\psi^{N}({\mathcal{P}}))\times\tau({\mathcal{P}}^{-1}))\right)$
$={}$ $p_{13*}\delta^{!}\left(\left({\mathrm{td}}(-q_{1}^{*}T_{\overline{J}_{C}}-q_{2}^{*}T_{\overline{J}_{C}})\cap\tau(\psi^{N}({\mathcal{P}}))\right)\times({\mathrm{td}}(-\pi_{2}^{*}T_{B})\cap\tau({\mathcal{P}}^{-1}))\right)$
$={}$ ${\mathfrak{F}}^{-1}\circ\widetilde{{\mathrm{ch}}}(\psi^{N}({\mathcal{P}})).$
The dimension bound of Corollary 3.5 now implies that both ends of (61) vanish in codimensions <g<g. In other words, we have for $k<2g$,
@@ 3.6. Variants
Recall from [5, 6] that Arinkin’s normalized Poincaré sheaf $\mathcal{P}$} on $\overline{J}_{C}\times_{B}\overline{J}_{C}$ is constructed from the normalized universal family $\mathcal{F}$} on $C\times_{B}\overline{J}_{C}$ of degree 0 rank $1$ torsion-free sheaves. Here the normalization of $\mathcal{F}$} is characterized by the trivializations along the sections
@@
Using the section of $C\to B$, we can identify the degree $d$ compactified Jacobian $\overline{J}^{d}_{C}$ with the (degree 0) compactified Jacobian $\overline{J}_{C}$. A universal family of rank $1$ degree $d$ torsion-free sheaves on $C\times_{B}\overline{J}^{d}_{C}=C\times_{B}\overline{J}_{C}$ is given by
@@ 4.2. Universal families and gerbes
This is representable by a Deligne–Mumford stack. Finally it is a $\mu_{r}$-gerbe over $\overline{J}^{d}_{C}$ since the perfect complex $p_{T*}({\mathcal{F}}^{d}_{T}|_{D\times_{B}T})$ has rank $r$ (which can be seen directly from nonsingular fibers). ∎
@@
It suffices to show that their cones in the derived categories are 0; this is a property which can be checked étale locally over the base $B$. We postpone the étale local calculations to Corollary 4.5 below. ∎
@@
They are independent and do not have any nontrivial relation. Due to the existence of the section, the relative compactified Jacobian $\overline{J}^{d}_{C}$ are identified for any choices of $d$. As before, we denote by $\overline{J}_{C}$ the degree 0 compactified Jacobian, and we want to compare ${\mathcal{F}}^{d}$ on $C\times_{U}\overline{{\mathcal{J}}}^{d}_{C}$ (which is trivialized along the multisection $D$) with the normalized universal sheaf $\mathcal{F}$} on $C\times_{U}\overline{J}_{C}$ (which is trivialized along the section $U$).
@@
Note that $\overline{J}_{C}$ is a fine moduli space. Therefore, describing a morphism from $\overline{{\mathcal{J}}}^{d}_{C}$ is equivalent to constructing a family of rank $1$ torsion-free sheaves of degree 0 trivialized along the section $U\subset C$, for which we may use ${\mathcal{F}}^{d}$. More precisely, we define
@@
Here the second factor adjusts the degree from $d$ to 0, and the third factor is included to trivialize the universal sheaf along the section. This yields a morphism $\sigma_{d}:\overline{{\mathcal{J}}}^{d}_{C}\to\overline{J}_{C}$ satisfying
@@
The normalized Poincaré sheaf $\mathcal{P}$} and its inverse ${\mathcal{P}}^{-1}$ are canonically defined on the target. The following is an immediate consequence of Proposition 4.4 as in Section 3.6 (using the argument of [1, Proposition 3.1(i)]). It provides an étale local description of the Poincaré sheaf ${\mathcal{P}}_{e,d}$ and its inverse ${\mathcal{P}}^{-1}_{e,d}$.
@@ 5.3.1. The locus of integral curves
$\mathrm{supp}(F)]=rH,\quad\chi(F)=\chi.$
@@
[…]
@@
This is well-known: the component given by the closure of reducible curves with two components of bidegrees (1,r1)(1,r-1) has the largest dimension among all the boundary components. ∎
@@
$i_{*}i^{!}{\mathbb{Q}}_{M_{r,\chi}}\to{\mathbb{Q}}_{M_{r,\chi}}\to j_{*}j^{*}{\mathbb{Q}}_{M_{r,\chi}}\to i_{*}i^{!}{\mathbb{Q}}_{M_{r,\chi}}[1]$
@@
(71) $h_{*}i_{*}i^{!}{\mathbb{Q}}_{M_{r,\chi}}\to h_{*}{\mathbb{Q}}_{M_{r,\chi}}\xrightarrow{~~(*)~~}j_{*}h_{W*}{\mathbb{Q}}_{M_{W}}\to h_{*}i_{*}i^{!}{\mathbb{Q}}_{M_{r,\chi}}[1].$
The filtered morphism $\mathrm{res}_{W}$ is induced by taking the global cohomology of ()(*). By Lemma 5.2, we have
$i_{*}i^{!}{\mathbb{Q}}_{M_{r,\chi}}\in D_{c}^{\geq 2(r-1)}(M_{r,\chi});$
in particular, both the first and the last terms of (71) are concentrated in degrees >2(r2)>2(r-2). Therefore the restriction map
$\mathrm{res}_{W}:H^{\leq 2(r-2)}(M_{r,\chi},{\mathbb{Q}})\to H^{\leq 2(r-2)}(M_{r,\chi},{\mathbb{Q}})$
is an isomorphism that preserves the perverse filtrations:
$\mathrm{res}_{W}\left(P_{k}H^{\leq 2(r-2)}(M_{r,\chi},{\mathbb{Q}})\right)\subset P_{k}H^{\leq 2(r-2)}(M_{r,\chi},{\mathbb{Q}}).$
Furthermore, this inclusion has to be an isomorphism due to [13, Lemma 3.3]. The corollary follows. ∎